OpenStudy (anonymous):

What is the area of a sector with a central angle of 160 degrees and a diameter of 5.8m

OpenStudy (anonymous):

\[\frac{160}{360} \pi \left(\frac{58}{20}\right)^2=\frac{841 \pi }{225}=11.7426 \]